# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v2

@@ Discussion
We introduced a novel method, CODERS, to identify the onset of careless responding (or absence thereof) for each participant in a given survey. It is motivated by literature stressing that in lengthy surveys—which are common in the behavioral and organizational sciences—a large proportion of survey participants may respond partially carelessly due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors as an indicator of inconsistency, and the proposed LongstringPattern (LSP) sequence as an indicator of invariability. Re-analyzing data from a seminal study on careless responding, we find evidence of partial careless responding that was hitherto undetected. Additionally, simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness.
@@
While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research. For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work. In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
@@
Whereas we apply machine learning and nonparametric techniques in our approach, other literature is focused on model-based approaches. For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding which might be violated in practice.
@@
Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding. All computations were performed with R version 4.3.3 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009). Our implementation of CODERS in the R package carelessonset is available from https://github.com/mwelz/carelessonset. Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
@@ Appendix A Setup and Assumptions
Throughout our paper, we consider the following setup. Let $\bm{X}$ be an $n\times p$ data matrix holding the rating-scale responses of $n$ survey participants to $p$ items. We make the standard assumption that all observations (i.e., all participants whose responses the survey collects) are independently and identically distributed. Moreover, it is a priori unknown if and when careless responding occurs in $\bm{X}$. We require the following assumptions.
@@
The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation, $\bm{S}$, where $s\ll p$. The dimension $s$ is known and corresponds to the number of constructs that the survey measures.
@@
The items in the survey are in a randomized order.
@@
All participants begin the survey as attentive respondents by providing accurate and truthful responses. As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
@@
The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors and LongstringPattern (LSP) sequences as indicators.
Assumption 1 is very mild in survey data because surveys typically measure multiple constructs and the number of constructs is typically known. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
Assumption 2 is again mild in survey data. There exist well-established and reliable survey measures for a large variety of variables. For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, the omega coefficient and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures). Hence, Assumption 2 is satisfied when the survey data were collected with measures of high reliability.
Assumption 3 is common practice in surveys in the social and behavioral sciences. The randomized order of items prevents that recurring response patters emerge as a result of the survey design rather than due to invariable careless responding.
Assumption 4 is discussed in the section on limitations, and Assumption 5 is motivated in detail in the section introducing the proposed methodology.
Appendix B Technical Details on CODERS
B.1 Details on Autoencoders
B.1.1 Network Architecture and Estimation
@@ B.1.2 Modeling Choices
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the presudo-Huber loss (B.2), we set constant $\delta=1$.
B.2 LongstringPattern Algorithm
@@ 7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
B.3 Details on Changepoint Detection

### v3

@@ Discussion
We introduced a novel method, CODERS, designed to identify the onset of careless responding (or absence thereof) for each participant in a survey. It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral and organizational sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors (RE) as an indicator of inconsistency, and the proposed LongStringPattern (LSP) sequence as an indicator of invariability. Simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness. Re-analyzing data from a seminal study on careless responding, we demonstrate how CODERS reveals novel insights on partial carelessness. In particular when combined with a post-hoc analysis of the identified response segments, CODERS is able to uncover various types of changes in (careless) response behavior.
@@
While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research. For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work. In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research. Furthermore, it could be interesting to study whether autoencoder reconstruction errors, which can be thought of comparing observed responses with an expectation, are able to distinguish between inconsistent careless responding and other sources of deviations from expected response behavior, such as misunderstanding or misinterpretation of survey items.
CODERS naturally integrates with existing techniques that address careless responding. First, as illustrated in our empirical application, traditional post-hoc screeners and CODERS provide complementary insights. The former aim at identifying which respondents are careless, whereas CODERS specifically looks for changepoints in response behavior. It should be noted that some implementations allow traditional screeners to be computed on subsets of responses. For instance, the implementation of the intra-individual variance (IRV) (Dunn et al.,, 2018; Marjanovic et al.,, 2015) in the R package careless (Yentes & Wilhelm,, 2023) can be calculated for a given number of subsets of equal length. Although this approach is related to the detection of periods of careless responding, it analyzes predefined subsets of responses, while CODERS aims to identify the location of a changepoint.
Second, pre-administered detection items such as bogus, instructed, or self-report items (e.g. Arthur et al.,, 2021; Curran,, 2016; Ward & Meade,, 2023) may provide further validation for the findings of CODERS. An advantage of detection items is that they can in principle detect any type of careless responding. By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade & Craig, (2012) to include at most three detection items, spaced every 50–100 items. However, detection items can strengthen the findings of CODERS: if passed detection items (i.e., responses expected from an attentive participant) fall with the found attentive period while failed detection items fall within the found careless period, there is convergent evidence of such response behavior.
Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson & Bowling,, 2020), or threatening with punishment for careless responding (Bowling et al.,, 2021). There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012). However, preventive measures can be undesirable or even unethical in practice. For instance, Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al., (2021) stress that promising rewards may cross ethical boundaries in organizational surveys. Notwithstanding, if preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
@@
Whereas we apply machine learning and nonparametric techniques in CODERS, other literature is focused on model-based approaches. For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding.
@@
Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding. All computations were performed with R version 4.4.2 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009). Our implementation of CODERS in the R package carelessonset (Welz & Alfons,, 2024) is available from https://github.com/mwelz/carelessonset. Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
@@ Appendix A Setup and Assumptions
Throughout our paper, we consider the following setup. Let $\bm{X}$ be an $n\times p$ data matrix holding the rating-scale responses of $n$ survey participants to $p$ items. We make the standard assumption that all observations (i.e., all participants whose responses the survey collects) are independently and identically distributed. Moreover, it is a priori unknown if and when careless responding occurs in $\bm{X}$. We require the following assumptions. Importantly, these assumptions should not be understood as a conceptual model for careless responding, but rather as statistical prerequisites for CODERS to work best. In the immortal words of John von Neumann, “[the truth] is much too complicated to allow anything but approximations.”
@@
The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation $\bm{S}$, where $s\ll p$. The dimension $s$ is known and corresponds to the number of constructs that the survey measures.
This assumption is very mild in survey data since surveys typically measure multiple constructs, with the number of constructs typically known to the survey designer. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
@@
This assumption is again mild in survey data. There exist well-established and reliable survey measures for a large variety of constructs. For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures). Hence, this assumption is satisfied when the survey data are collected with measures of high reliability.
@@
The survey design precludes responses in (fixed) recurring patterns, meaning that such patterns are indicative of invariable careless responding.
One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the social and behavioral sciences. If it is desirable to keep items of the same construct together to reduce the burden on respondents, as is often the case in organizational surveys, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together). The use of negatively keyed items may also be beneficial.
@@
Participants begin the survey as attentive respondents by providing accurate and truthful responses. As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
Under this assumption, a detected changepoint (see also the next assumption) naturally divides a participant’s responses into an attentive segment before the changepoint and a careless segment from the changepoint onward. In practice, a changepoint may characterize different changes in response behavior. For instance, a participant may engage in careless responding in the beginning of the survey but later switch to attentive responding, or they may switch from one type of careless responding to another (e.g., from inconsistency to invariability). Via a post-hoc analysis of the identified response segments, CODERS remains useful for uncovering insights on such changes in response behavior, as highlighted in the empirical application with selected cases (cf. Figures 11(c)11(e) and Tables 3(c)3(e) in the main text).
It is also conceivable that a respondent may switch repeatedly between attentive and careless responding. For instance, they may start attentively, then be distracted for a certain period of time, but switch back to attentive responding once the distraction goes away. CODERS in its current form is not designed to pick up such behavior, but an extension using multiple changepoint detection may be a fruitful venue for future research.
@@
The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors (RE) and LongStringPattern (LSP) sequences as indicators.
A detailed motivation for this assumption is provided in the section introducing CODERS.
Appendix B Details on CODERS
B.1 Motivating Novel Methodology
Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as psychometric antonym, personal reliability, and the longstring index (see e.g., Alfons & Welz,, 2024, for a recent overview of traditional screeners).
A primary motivation for devising new methodology is that most existing screeners compute a single carelessness score for each participant. While this can work very well for distinguishing careless from attentive respondents, having a singular score for every respondent does not provide much information beyond the attentive-careless classification. As outlined in the main text, recent literature suggests that partially careless responding might be a common occurrence. As such, it might be more informative to know when a given respondent starts responding carelessly (if at all) rather than a dichotomous attentive-careless distinction. For instance, this could be useful for knowing when the responses of a given respondent become unreliable so that one must not drop all of their responses in a primary analysis. In order to obtain a granular estimate of carelessness onset on the participant level, we require a carelessness score for every item for every respondent rather than a singular per-participant carelessness score. While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al.,, 2018; Marjanovic et al.,, 2015), accurate estimation of the onset of carelessness necessitates new methodology.
As such, CODERS is based on two different carelessness scores on the participant-item level to capture a wide variety of different types of careless responding. The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level. Autoencoders are described in detail in Section B.2. The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson,, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $-2-1-2-\dots-1-$ 2 - ⋯ - 1 - 2. If so, it allows one to trace back when the recurring sequence has started, thereby providing evidence for the onset of invariable carelessness. LSP is explained in detail in Section B.3.
B.2 Details on Autoencoders
B.2.1 Further Intuition
Assigning a participant-item score that measures how inconsistently a participant has responded to the item in question arguably is a challenging task. Since different attentive participants may respond differently to a given item (between-participant heterogeneity), one needs to somehow understand how attentive responding manifests on the participant-level so that a possible switch to inconsistent responding can be detected. Otherwise, there would be the risk of confusing attentive responding for inconsistent responding, and vice versa. We argue that an auto-associative neural network (autoencoder) is useful for learning within-participant attentive response structures.
An autoencoder as used in CODERS essentially predicts the response of each respondent to each item. The underlying idea is that if the predicted response is notably different from the given response, the participant has given a response that is unexpected, judging from their previous response behavior. While an occasional unexpected response is not out of the ordinary, having multiple such unexpected responses occur consecutively may suggest that the participant has shifted from predictable responding to unpredictable responding, which might be viewed as indication of an onset of inconsistent responding. Since we assume that every participant has provided attentive responses at the beginning of the questionnaire (see Assumption 4), it should be conceptually possible to learn how attentive responding manifests for each participant so that somewhat accurate response predictions can be made up to carelessness onset.
The autoencoder learns attentive response behavior by exploiting a key feature of questionnaires, namely that multiple items measure the same construct through psychometric scales. Since one usually knows how many scales are measured by the survey at hand, the information contained in a given participant’s responses should be representable in much fewer dimensions than the total number of items (see Assumption 1). Thus, the task of learning attentive response behavior can be seen as finding a good lower-dimensional representation for every participant’s given responses.
Formally, the autoencoder minimizes the participant-item response prediction error over all participants and items. However, in the process of making predictions, the autoencoder enforces to first find a lower-dimensional representation of each participant’s given responses, and then to make a response prediction based on that lower-dimensional representation. In other words, the autoencoder first needs to compress $p$ dimensions into $s\ll p$ dimensions, and then decompress again that lower-dimensional representation to obtain $p$-dimensional predictions. The potential presence of inconsistent careless responses should not much affect the quality of the lower-dimensional representation—unless carelessness onsets very early—because inconsistent responding is characterized by its random-like nature, meaning that such responses are not informative for learning the lower-dimensional representation.
We stress that response predictions based on a lower-dimensional representation could also be obtained through principal component analysis (PCA). In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer,, 1991). Since we wish to refrain from possibly restrictive assumptions, we choose autoencoders over PCA to construct lower-dimensional representations of the given responses. We are now ready to dive into the formal definition of the autoencoder.
B.2.2 Network Architecture and Estimation
@@ B.2.3 Modeling Choices
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the pseudo-Huber loss (B.2), we set constant $\delta=1$.
B.3 LongStringPattern Algorithm
@@ 7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
B.4 Details on Changepoint Detection